Kagiso wants to buy a new gaming computer for R40 000. He decides to save by depositing an amount of R400 quarterly into an account earning 16% interest per year, compounded quarterly. The approximate number of quarters it will take Kagiso to have R40 000 available is


1
Expert's answer
2022-03-16T17:52:41-0400

Solution


Interest rate = 16% per year = 16/12 % per month = 4/3 % per month


Future Value FV = R 40,000


Deposit Per month A = R 400


Number of periods required to generate R 40,000 is n =?


"\begin{gathered}
 n = \ln \left( {1 + \frac{{\left( {FV \times r} \right)}}{A}} \right) \div \ln \left( {1 + r} \right) \\
 n = \ln \left( {1 + \frac{{\left( {40,000 \times \frac{1}{{75}}} \right)}}{{400}}} \right) \div \ln \left( {1 + \frac{1}{{75}}} \right) \\
 n = 63.97 \\ 
\end{gathered}"


Hence it will Kagiso to take around 64 months to get the amount R 40,000


The same has been displayed by the following table